OpenStudy (anonymous):

MEDAL FOR BEST ANSWER Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

divide -3 on both sides then tell me wat u get

OpenStudy (anonymous):

-11?

OpenStudy (anonymous):

-3|x + 7| = -12 \[|x + 7| = \frac{-12}{-3} \rightarrow |x + 7| = 4\] \[ x + 7 = \pm 4\] i.e. \[ x + 7 = 4\] or x + 7 = - 4 i.e. x= 4-7 or x= -4-7 i.e. x= -3 or x= -11 is the required answer. @lilman212

OpenStudy (anonymous):

no i just said to divide by -3 on both sides not to solve the whole problem

OpenStudy (anonymous):

i wanted to see if u knew how to work the problem just not to give u the answer

OpenStudy (anonymous):

we can simply do -3|x + 7| = -12 |-3x-21|=-12 3x+21=-12 asafter retriving any value from modulas the sign is changes to positive x=-12-21/3 x=-11